A speeding car passes a highway patrol checkpoint and then accelerates at a constant rate. 3.15 s later, the car is 240 m from the checkpoint and its speed is then 33.7 m/s. What is the acceleration of the car?
1
Expert's answer
2017-02-23T06:05:40-0500
[s = 240 m ]
[V1 - speed at the checkpoint ]
[V2 = 33.7 m/s]
[a - acceleration]
[t = 3.15 s]
s = V1*t + (at^2)/2 (1)
V2 = V1 + at => V1 = V2 - at (2)
(2) -> (1) :
s = (V2-at)t + (at^2)/2 = V2*t - at^2 + (at^2)/2 = V2*t - (at^2)/2 (3)
From (3):
a = 2*(V2*t - s)/(t^2)
a = -27 [ m/(s^2) ]
So, V1 = 118.7 (m/s)

Answer: a = -27 [ m/(s^2) ]